OpenStudy (anonymous):

graph the equation using the point and the slope. y+9=1over2(x-2)

OpenStudy (anonymous):

y - (-9) = 1/2 ( x - (2) so your point is (2,-9) y = 1/2x - 10

OpenStudy (anonymous):

thanks lol i have another one coming at you.